# A silicon spin vacuum: isotopically enriched 28silicon-on-insulator and 28silicon from ultra-high fluence ion implantation

### v1

@@ III.4 Open volume defects
Figure 4: TEM images of 28Si enriched samples on bulk Si where the gold decoration (diffusion) experiment was performed (a) after and (b) during the SPE anneal stage. The thin carbon protective layer was deposited onto the sample surface during TEM sample preparation. Gold surface segregation and decoration of voids are indicated by red arrows. A higher magnification TEM image of the region highlighted by the dashed white box is shown as an inset figure. (c) RBS measured gold depth profile for a pristine Si sample after Au diffusion and the 28Si enriched samples shown in the TEM images (a) and (b). The vertical axis (normalized counts) is proportional to the Au concentration.
@@
Our experiments indicate that the observation of open volume defects (vacancies or voids) through gold decoration depends crucially on at which stage the gold is introduced. In Fig. 4(a) and (b), we show TEM results from the 28Si enriched bulk Si sample (60 keV $0.77\times 10^{18}$ cm-2, high flux, RT) where the gold decoration (diffusion anneal) was performed either after or during the SPE annealing stage. Specifically, the sample in Fig. 4(a) was prepared by first performing the SPE anneal (after ion implantation) followed by gold deposition and subsequent annealing to in-diffuse the surface gold. On the other hand, the sample in Fig. 4(b) was prepared by first depositing gold on the sample surface (after ion implantation) followed by SPE annealing. In this case, the gold in-diffusion occurs at the same time as the SPE recrystallization.
@@
In the sample where the gold is diffused into the sample after SPE annealing [Fig. 4(a)], TEM shows a fully recrystallized Si layer (supported by SAED measurements, not shown) with no evidence of gold-decorated open volumes. Rather, the gold appears as a thin dark layer segregated at the surface underneath the carbon protective layer. This is confirmed by RBS profiling of the gold that we will discuss later. On the contrary, in the sample where the gold is in-diffused during the SPE annealing stage [Fig. 4(b)], the open volume defects (voids) can be clearly observed up to a depth of 70 nm from the sample surface. They appear as circular regions of darker contrast in the TEM images due to the high mass contrast of gold (confirmed by energy dispersive X-ray spectroscopy mapping, data not shown) and the gold atoms preferentially clustering within these voids. [35] Two such examples are indicated by the red arrows. The diameter of these voids range from 1 nm to 7 nm. Both the size and density distribution of these gold-decorated voids achieve a maximum at/near the surface and decrease deeper into the sample. These voids are no longer visible under TEM at depth regions beyond 70 nm. Additionally, we also observe stacking-like defects at the surface. An example of this is highlighted by the white dashed region in Fig. 4(b). A higher magnification TEM image of this region is shown in the inset image. These stacking-like defects were not observed in our other samples that were SPE annealed in the absence of gold at the surface [see, for example, Fig. 1(c) and Fig. 4(a)]. Thus, we attribute these stacking defects to the distortion of the SPE planar growth front due to local modifications to the SPE rate in the presence of gold. [55, 23] The dark band of defects located at 200 nm from the surface are the EOR defects, as observed and discussed previously [Fig. 1]. We note that the TEM images shown in Fig. 4(a) and (b) were taken under the same magnification settings and the slight difference (7 nm) in the EOR depth location observed between them can be accounted for by TEM scale calibration errors (5%).
@@
RBS-C measurements of the gold depth profiles in the same gold-decorated 28Si samples shown in Fig. 4(a) and (b) are shown in Fig. 4(c). A reference gold profile obtained from an un-implanted sample (‘Pristine’) diffused with gold under the same annealing conditions (620C for 10 min) is also shown. The vertical axis (‘Normalized counts’) is proportional to the gold concentration. In the pristine case, the gold remains predominantly at the surface with a diffusion tail that is 100 times smaller in magnitude that is discernible between the surface and a depth region of 100 nm. This is expected since gold is a fast diffuser and is known to preferentially segregate at the front and back surfaces of Si. [56] We note that the gold surface peak appears slightly broadened at the surface (depth = 0 nm) due to the energy resolution of our RBS detector. On the other hand, the gold distribution for the gold-decorated 28Si samples depends on when the Au diffusion annealing was performed. If the gold diffusion is performed after the SPE annealing step, we observed no significant difference between the pristine sample and 28Si gold profiles (solid black circles and red stars, respectively), except a small gold peak that is observed in the 28Si sample at a depth location greater than 200 nm caused by gold gettering at the EOR defects. The EOR gold gettering peak was observed in both 28Si samples irrespective of when the gold diffusion anneal was performed and is indicated by the arrows and text in Fig. 4(c). The depth location of the EOR gold gettering observed in RBS ($210\pm 20$ nm) is in good agreement with TEM ($220\pm 10$ nm).
@@
If the gold diffusion and SPE annealing are performed simultaneously, we observe in RBS significant accumulation of gold between the surface and 50 nm as compared to the pristine sample, with a concentration tail that decreases and extends deeper into the Si layer. Unsurprisingly, the depth region where the gold is accumulating the most (0 to $50\pm 10$ nm) overlaps completely with the TEM depth region where the nanoscopic gold-decorated voids were observed (0 to $70\pm 10$ nm) in Fig. 4(b). Thus, the RBS gold distribution in this sample also reflects the depth profile of the gold-decorated voids. Both of these RBS and TEM observed depth regions are also in good agreement with the calculated depth region for excess vacancies, that is, between the surface and $R_{p}/2$. For the samples shown in Fig. 4, $R_{p}/2=44$ nm and $\Delta z=15$ nm calculated from SRIM and TRIDYN, respectively, and thus excess vacancies are expected up to a depth of 60 nm.
Our experiments above suggest that the nanoscopic voids observed in Fig. 4(b) were formed from vacancy excesses during implantation and fully stabilized from gold-vacancy interactions in amorphous Si following 28Si ion implantation. In the sample where the SPE annealing and gold diffusion occurred simultaneously [Fig. 4(b)], the direction of the SPE recrystallization growth front (towards the surface) and gold diffusion (away from the surface) are directly opposite to each other. At an annealing temperature of 620C, the SPE growth rate and the diffusion coefficient of gold in amorphous Si are estimated to be 2 nm/s and 1 nm2/s, respectively. [57, 58]. For an a/c-interface (EOR) that is 220 nm deep, the SPE growth front and gold diffusion tail are therefore expected to meet at a depth location of about 48 nm from the surface after 86 s of annealing. Both the further diffusion of gold in crystalline Si beyond the SPE growth front and EOR (towards the Si back surface) and the complete recrystallization of the amorphous Si layer subsequently occur over the remaining annealing duration (about 9 minutes). The competition between gold diffusion and SPE growth and their meeting at around 48 nm may therefore explain the observed abrupt decrease in the gold concentration at 50 nm and the size distribution of the gold-decorated voids in our sample (‘${}^{28}Si$ gold during SPE’).
To obtain a better understanding of the physical origin of these open volumes (vacancies and voids), we performed additional gold decoration experiments on selected 28SOI samples implanted under different temperature and flux conditions where the gold was introduced during the SPE annealing stage. The experimental data for these samples are shown in the Supplementary Material but the key results will be described here. RBS gold profiling indicates that gold had accumulated between the surface and about 50 nm in all of our implanted 28SOI samples. The gold distribution in these samples also appears to depend on the ion flux and sample temperature. Higher fluxes and lower temperatures during implantation result in a higher total gold concentration accumulating below the surface. Indeed, the excess vacancy concentration in the amorphous Si phase is expected to depend on both the ion flux (which dictates the point defect production and clustering rates) and sample temperature (which controls the defect annihilation and clustering rates). Higher stable vacancy excesses due to high ion flux and/or lower sample temperatures may then promote subsequent formation of voids that are revealed by gold decoration during gold diffusion in amorphous Si. [58] Whether the voids are fully formed from vacancy excesses in amorphous Si during implantation, or vacancy/void clustering is enabled by gold interactions is not known. However, we note that even if these open volumes were formed during the ion implantation process, they can be completely removed by a 620C 10 minute anneal and are only stabilized by the presence of external impurities such as gold. Further experiments will be required to determine the exact physical origin of these open volumes.
Finally, we note that it may also be possible for the nanoscopic voids and/or high vacancy excesses to cause some degree of surface swelling and a more positive $\Delta z$ value. Void swelling, as it is termed in the literature, is unfortunately difficult to quantify and is rarely addressed in the published literature. [59, 60, 61] Although these voids can be readily removed by SPE annealing, it may partially account for the observed larger $\Delta z$ values and the slightly larger surface binding energies required to fit the simulated curve to our experiments.
@@ IV Concluding remarks
Our experiments and simulations demonstrate the ability of the 28Si enrichment process to precisely control the purity of 28Si layers formed on bulk Si and SOI substrates. We note that it was necessary to increase the surface binding energy by 10% to a value of 5.17 eV in TRIDYN to obtain a better agreement with experimentally observed surface sputtering and 29Si depletion characteristics. Key requirements to produce 28Si surface layers with residual 29Si and 30Si concentrations below 1 ppm and at least 100 nm thick include choosing ion implantation energies that result in surface accumulation ($<3$ keV or $>40$ keV) and ion fluences exceeding $0.6-0.9\times 10^{19}$ cm-2. We produce one such sample with 60 keV 28Si ions implanted in bulk Si up to a fluence of $0.77\times 10^{19}$ cm-2, where the 29Si and 30Si isotopes were depleted to less than 7 ppm, a value limited by our SIMS resolution. Further, our experiments emphasize the importance of performing 28Si enrichment experiments under high vacuum ($<5\times 10^{-8}$ Torr) and high ion flux conditions to minimize contamination from background residual gas atoms. High ion fluxes also reduces the overall implantation duration and is desirable for producing samples with larger areas for scaling up. For SOI substrates, high flux ion implantation above $70\ \mu$A/cm2 performed at RT results in significant broadening of the a/c-interface and interstitial clustering which, after subsequent SPE annealing, resulted in the formation of threading dislocations throughout the 28Si layer. This was not observed in bulk Si substrates and can be circumvented by performing the ion implantation in SOI at 77 K to suppress dynamic annealing. We also report the direct observation of gold-decorated voids in our 28Si samples. The observation of these voids depend critically on when the gold diffusion anneal is performed. In this work, they were only observed when gold is introduced into the amorphous layer during SPE annealing before the growth front catches up to the gold diffusion tail. These voids are no longer visible if the gold atoms are introduced after complete SPE recrystallization of the amorphous implanted layer, which is performed under moderate temperature annealing at 620C for 10 minutes. Future low energy positron annihilation spectroscopy experiments may reveal further information related to the excess vacancy distribution in our samples. From a Si device technology and QIP application perspective, our work provides a clear pathway towards realizing large-scale 28Si and 28SOI spin vacuum substrates for ultra-long qubit coherence applications. Spin resonance and qubit spectroscopy experiments are currently underway to uncover the coherent lifetimes of donor qubits embedded in our 28Si and 28SOI substrates.

### v2

@@ III.4 Open volume defects
Figure 4: TEM images of 28Si enriched samples on bulk Si where the gold decoration (diffusion) experiment was performed (a) after and (b) during the SPE anneal stage. The thin carbon protective layer was deposited onto the sample surface during TEM sample preparation. Gold surface segregation and precipitation are indicated by red arrows. A higher magnification TEM image of the region highlighted by the dashed white box is shown as an inset figure. (c) RBS measured gold depth profile for a pristine Si sample after Au diffusion and the 28Si enriched samples shown in the TEM images (a) and (b). The vertical axis (normalized counts) is proportional to the Au concentration.
@@
Figure 4(a) and (b) show TEM results from the 28Si enriched bulk Si sample (60 keV $0.77\times 10^{18}$ cm-2, high flux, RT) where the gold in-diffusion anneal was performed either after or during the SPE annealing stage, respectively. Specifically, the sample in Fig. 4(a) was prepared by first performing the SPE anneal (after ion implantation) followed by gold deposition and subsequent annealing to in-diffuse the surface gold. On the other hand, the sample in Fig. 4(b) was prepared by first depositing gold on the sample surface (after ion implantation) followed by SPE annealing. In this case, the gold in-diffusion occurs at the same time as the SPE recrystallization. In both Figs. 4(a) and (b), the dark band of defects located at 200 nm from the surface are the EOR defects, as observed and discussed previously [Fig. 1]. Both of these TEM images were taken under the same magnification settings and the slight difference (7 nm) in the EOR depth location observed between them can be accounted for by TEM scale calibration errors (5%).
@@
In the sample where the gold is in-diffused after SPE annealing [Fig. 4(a)], TEM shows a perfectly recrystallized Si layer (supported by SAED measurements, not shown). The gold atoms appear as a thin dark layer segregated at the surface underneath the carbon protective layer. This is confirmed by RBS profiling of the gold that we will discuss later. On the contrary, in the sample where the gold is in-diffused during the SPE annealing stage [Fig. 4(b)], circular regions of darker contrast can be clearly observed up to a depth of 70 nm from the sample surface. These are regions that contain higher concentrations of gold (confirmed by energy dispersive X-ray spectroscopy mapping, data not shown) and are darkened due to the high mass contrast of gold. Two such examples are indicated by the red arrows. We will tentatively refer to them as ‘gold-rich blobs’ and discuss their significance in relation to open volume defects later. The diameter of these gold-rich blobs range from 1 nm to 7 nm, and their size and density distribution achieve a maximum at/near the surface and decrease deeper into the sample. They are no longer visible under TEM at depth regions beyond 70 nm.
Additionally, we observe stacking-like defects at the surface of the sample in Fig. 4(b). An example of this is highlighted by the white dashed region. A higher magnification TEM image of this region is shown in the inset image. These stacking-like defects were not observed in our other samples that were SPE annealed in the absence of gold at the surface [see, for example, Fig. 1(c) and Fig. 4(a)]. Thus, we attribute these stacking defects to the distortion of the SPE planar growth front due to local modifications to the SPE rate in the presence of gold. [56]
@@
RBS-C measured gold depth profiles of the gold diffused samples are shown in Fig. 4(c). A reference gold profile obtained from an un-implanted sample (‘Pristine’) in-diffused with gold under the same annealing conditions (620C for 10 min) is also shown. The vertical axis (‘Normalized counts’) is proportional to the gold concentration. In the pristine case, the gold remains predominantly at the surface with a diffusion tail that is 100 times smaller in magnitude that is discernible between the surface and a depth region of 100 nm. We note that the gold surface peak appears slightly broadened at the surface (depth = 0 nm) due to the energy resolution of our RBS detector.
@@
On the other hand, the gold distribution in the 28Si samples depends crucially on when the gold in-diffusion was performed. If the gold in-diffusion is performed after the SPE annealing step, we observed no significant difference between the pristine Si and 28Si gold profiles (solid black circles and red stars, respectively), except a small gold peak that is observed in the 28Si sample at a depth location greater than 200 nm caused by gold gettering at the EOR defects. The EOR gold gettering peak was observed in both 28Si samples irrespective of when the gold in-diffusion was performed and is indicated by the arrows and text in Fig. 4(c). The depth locations of the EOR gold gettering peaks observed in RBS ($210\pm 20$ nm) are in good agreement with TEM ($220\pm 10$ nm). If the gold in-diffusion and SPE annealing are performed simultaneously, we observe in RBS significant accumulation of gold between the surface and 50 nm as compared to the pristine sample, with a concentration tail that decreases and extends deeper into the Si layer. Unsurprisingly, the depth region where the gold is accumulating the most (0 to $50\pm 10$ nm) overlaps completely with the TEM depth region where the nanoscopic gold-rich blobs were observed (0 to $70\pm 10$ nm).
We will now discuss the physical origins of these gold-rich blobs and their sub-surface depth distribution. In our pristine Si sample and 28Si sample where the gold in-diffusion was performed after SPE annealing, the in-diffusion of gold occurs entirely in the Si crystalline phase. In crystalline Si, gold diffuses via the kick-out mechanism, where an interstitial gold atom substitutes a Si atom on a lattice site, thereby forming a Si self-interstitial. This diffusion mechanism enables gold to diffuse rapidly in Si, and, combined with a low gold solubility in crystalline Si, results in a U-shaped depth distribution with gold strongly segregating at the front and back surfaces of Si. [57] Indeed, this mechanism adequately explains the gold surface segregation observed in both our pristine and 28Si samples. However, in the 28Si sample, the perceived gold solubility in the vicinity of the EOR defects (dislocated Si) is higher than in the surrounding (dislocation-free crystalline) Si matrix due to the efficient gold gettering by dislocations at the EOR. [58]
On the other hand, in the 28Si sample where the gold was in-diffused during the SPE annealing, the initial in-diffusion of gold occurs in the amorphous Si phase. At an annealing temperature of 620C, the diffusion coefficient of gold in amorphous Si and the SPE growth rate are estimated to be 1 nm2/s and 2 nm/s, respectively. [59, 60] For an initial a/c-interface that is 220 nm deep, the gold in-diffusion front (moving into the substrate) and the SPE growth front (moving towards the surface) will meet at a depth location of 48 nm from the surface after 86 s of annealing. The nature of the gold atoms within the amorphous layer at this stage of the annealing process requires some discussion. Previous works by Elliman et al. [61] have shown that gold will be uniformly distributed within amorphous Si even at annealing temperatures as low as 450C. Gold precipitation was also observed within the amorphous Si layer following its diffusion, and, curiously, the gold precipitates were localized to only the upper half of the amorphous layer. Indeed, we believe that the gold-rich blobs in Fig. 4(b) are also gold precipitates that were formed in the amorphous Si layer prior to the completion of the SPE. However, rather than their dissolution and zone refinement, [56] the gold precipitates in our sample remain stable in the crystalline Si phase upon further SPE annealing. We note that zone refinement of gold in amorphous Si is typically expected for soluble gold in its unprecipitated form. It remains unclear whether the gold precipitates in our sample were formed in amorphous Si by decoration/generation of open volumes in the amorphous phase or were formed at the gold-saturated, amorphous-crystalline interface during SPE as a consequence of trapping of some of the segregating gold in the crystalline phase. Further work is needed to elucidate these proposals.
Finally, it is worth noting that in cases where initially crystalline Si is saturated with soluble gold (unprecipitated) through high temperature annealing, subsequent annealing at lower temperatures results in gold precipitation in the crystalline phase as the system evolves to thermal equilibrium (lower gold solubility) at that temperature. [62] During this process gold could diffuse via the dissociative mechanism [63] and precipitate at pre-existing defects or form a series of extrinsic stacking faults. In contrast, gold precipitation in amorphous Si as proposed in this work may be enabled by the presence of high concentrations of gold (and Si dangling bonds) within amorphous Si and thus the average diffusion length of gold in amorphous Si is shorter than the average separation between gold atoms.
@@ IV Concluding remarks
Our experiments and simulations demonstrate the ability of the 28Si enrichment process to precisely control the purity of 28Si layers formed on bulk Si and SOI substrates. We note that it was necessary to increase the surface binding energy by 10% to a value of 5.17 eV in TRIDYN to obtain a better agreement with experimentally observed surface sputtering and 29Si depletion characteristics. Key requirements to produce 28Si surface layers with residual 29Si and 30Si concentrations below 1 ppm and at least 100 nm thick include choosing ion implantation energies that result in surface accumulation ($<3$ keV or $>40$ keV) and ion fluences exceeding $0.6-0.9\times 10^{19}$ cm-2. We produced one such sample with 60 keV 28Si ions implanted in bulk Si up to a fluence of $0.77\times 10^{19}$ cm-2, where the 29Si and 30Si isotopes were depleted to less than 7 ppm, a value limited by our SIMS resolution. Further, our experiments emphasize the importance of performing 28Si enrichment experiments under high vacuum ($<5\times 10^{-8}$ Torr) and high ion flux conditions to minimize contamination from background residual gas atoms. High ion fluxes also reduces the overall implantation duration and is desirable for producing samples with larger areas for scaling up. For SOI substrates, high flux ion implantation above $70\ \mu$A/cm2 performed at RT results in significant broadening of the a/c-interface and interstitial clustering which, after subsequent SPE annealing, resulted in the formation of threading dislocations throughout the 28Si layer. This was not observed in bulk Si substrates and can be circumvented in SOI by performing the ion implantation at 77 K to suppress dynamic annealing. We were unable to detect (with TEM and gold decoration techniques) open volume defects in our 28Si samples after complete recrystallization by SPE annealing. Future low energy positron annihilation spectroscopy experiments may be better suited at measuring excess vacancies in our samples.
From a Si device technology and QIP application perspective, our work provides a clear pathway towards realizing large-scale 28Si and 28SOI spin vacuum substrates for ultra-long qubit coherence applications. Spin resonance and qubit spectroscopy experiments are currently underway to uncover the coherent lifetimes of donor qubits embedded in our 28Si and 28SOI substrates.